### v1

# Discrete Layered Entropy, Conditional Compression and a Tighter Strong Functional Representation Lemma

Abstract
We study a quantity called discrete layered entropy, which approximates the Shannon entropy within a logarithmic gap. Compared to the Shannon entropy, the discrete layered entropy is piecewise linear, approximates the expected length of the optimal one-to-one non-prefix-free encoding, and satisfies an elegant conditioning property. These properties make it useful for approximating the Shannon entropy in linear programming, studying the optimal length of conditional encoding, and bounding the entropy of monotonic mixture distributions. In particular, it can give a bound for the strong functional representation lemma that improves upon the best bound (as long as the mutual information is at least 2).
I Introduction
In this paper, we study a quantity which we call discrete layered entropy
$\Lambda(p):=\sum_{i=1}^{\infty}p^{\downarrow}(i)\left(i\log i-(i-1)\log(i-1)\right),$
where $p$ is a probability mass function, $p^{\downarrow}(i)$ is the $i$-th entry of $(p(x))_{x\in\mathcal{X}}$ when sorted in descending order. This is a discrete analogue of the (continuous) layered entropy studied in [1, 2]. Compared to the Shannon entropy $H(X)$, $\Lambda(X)$ has the following properties:
Figure 1: Left: Contour plot of the Shannon entropy $H(p)$ for $p:\{1,2,3\}\to[0,1]$ being a ternary probability mass function. Right: Contour plot of the discrete layered entropy $\Lambda(p)$. We can see that $\Lambda(p)$ is piecewise linear and lower-bounds $H(p)$.
  • $H(X)$ transforms nicer for joint random variable (e.g., $H(X,Y)=H(X)+H(Y)$ for independent $X,Y$), whereas $\Lambda(X)$ transforms nicer for conditioning ($\Lambda(X|Y)=\min_{U:H(X|Y,U)=0}\Lambda(U)$).
  • $H(X)$ approximates the expected length of the optimal prefix-free encoding of $X$, whereas $\Lambda(X)$ approximates the expected length of the optimal one-to-one non-prefix-free encoding [3, 4, 5] of $X$.
  • $\Lambda(X)$ is close to $H(X)$ within a logarithmic gap.
We present several applications of $\Lambda(X)$. First, it can be used to show a tighter strong functional representation lemma: for every $X,Y$, there exists $S$ independent of $X$ with $H(Y|X,S)=0$ and $H(Y|S)\leq I(X;Y)+\log(I(X;Y)+5.51)+1.06$, which improves upon previous results [6, 7, 8] for $I(X;Y)\geq 2$. We will also show another bound that improves upon [8] for $I(X;Y)\geq 0.7$. Second, it is a piecewise linear approximation of $H(X)$ useful for linear programming. Third, it is useful for conditional encoding tasks, to be described as follows.
Suppose the encoder wants to encode $X$, given a side information $Y$ that is also known to the decoder (i.e., the encoder encodes $X$ conditionally given $Y$). One method is to have the encoder find the ranking $U\in\mathbb{N}$ of $X$ among $p(x|Y)$ for all $x$ (i.e., $X$ has the $U$-th largest $p(x|Y)$), and sends $U$ to the decoder. This $U$ can be shown to be smallest possible compression, in the sense that it minimizes $H(U)$ (which is approximately the compression size using a prefix-free code) [^1] If we are allowed to encode $U$ conditional on $Y$ using a conditional prefix-free code, then the compression size would be $\approx H(U|Y)$. Nevertheless, if we require synchronization even when the decoder does not know $Y$, then we must use an unconditional prefix-free code, which has a length $\approx H(U)$. Refer to Section IV-B for discussions. subject to the recoverability requirement $H(X|Y,U)=0$. Hence, this $U$ can be regarded as the “conditional compression” of $X$ given $Y$, which contains the information in $X$ that is not in $Y$, and hence we use the notation $U=X\backslash Y$ to highlight the analogy to set difference.
It is perhaps unsatisfying that the entropy of the conditional compression $H(X\backslash Y)$ is generally not the conditional entropy $H(X|Y)$ (we only have $H(X\backslash Y)\geq H(X|Y)$). The conditional entropy $H(X|Y)$ does not actually correspond to the entropy of any random variable—it cannot be interpreted as the entropy $H(\cdot)$ applied to a “conditional random variable $X|Y$”. This is in stark contrast to the joint entropy $H(X,Y)$, which is indeed $H(\cdot)$ applied to the joint random variable $(X,Y)$.
We show that discrete layered entropy satisfies
$\Lambda(X\backslash Y)=\Lambda(X|Y),$ (1)
where $\Lambda(X|Y):=\mathbb{E}_{Y}[\Lambda(p_{X|Y}(\cdot|Y))]$ is defined in a similar manner as $H(X|Y)$. This fact is not only aesthetically pleasing (we can indeed treat “$X|Y$” as a random variable given by $X\backslash Y$), but also has several implications. First, this allows us to approximate the optimal expected encoding length of the aforementioned conditional encoding task by $\Lambda(X|Y)$. Second, it makes $\Lambda(X)$ a useful tool for bounding the entropy of a mixture of distributions with monotonic probability mass functions, since (1) is equivalent to saying that $\Lambda(X)$ is a linear function of the probability mass function $p_{X}$ when we restrict $p_{X}:\mathbb{N}\to\mathbb{R}$ to a nondecreasing function. We will utilize this to prove a tighter bound for the strong functional representation lemma [6] in Section VI.
Moreover, we show that (1) is the defining property of $\Lambda$, in the sense that the discrete layered entropy is the only function satisfying (1) and $\Lambda(X)=\log k$ for $X\sim\mathrm{Unif}(\{1,\ldots,k\})$. Also, $\Lambda$ is the largest function satisfying (1) and $\Lambda(X)\leq H(X)$, and hence $\Lambda(X)$ is the “best underestimate” of $H(X)$ that satisfies (1). We show that $\Lambda(X)$ is close to $H(X)$, in the sense that
$\Lambda(X)\leq H(X)\leq\Lambda(X)+\log\left(1+\frac{\Lambda(X)}{e\log e}\right)+\log e.$ (2)
This allows us to convert between $H(X)$ and $\Lambda(X)$ depending on whether we need the properties of $H$ or $\Lambda$ in the proof, incurring only a logarithmic gap for each conversion. This makes $\Lambda(X)$ a powerful tool even if we only want our final result to be in terms of Shannon entropy. For example, for the conditional encoding task, we can approximate the optimal encoding length $H(X\backslash Y)$ using (1) and (2) via $H(X\backslash Y)\approx\Lambda(X\backslash Y)=\Lambda(X|Y)\approx H(X|Y)$.
Related Works on Differences between Information
Apart from the “$X\backslash Y$” (which minimizes $H(U)$ subject to $H(X|Y,U)=0$) studied in this paper, there are several notions of differences between the information in $X$ and the information in $Y$ studied in the literature.
Strong functional representation lemma [6] seeks to minimize $H(X|U)$ subject to $U\perp\!\!\!\perp Y$ and $H(X|Y,U)=0$. These constraints are analogous to set differences $A\backslash B$ between sets $A,B$, which satisfies $(A\backslash B)\cap B=\emptyset$ and $A\subseteq(B\cup(A\backslash B))$. Note that $H(U)$ can be much larger than $H(X|Y)$. To review various bounds on $H(X|U)$, the result in [9] implies that $H(X|U)\leq I+(1+o(1))\log(I+1)$ can be achieved, where $I:=I(X;Y)$. [10] improved the bound to $I+\log(I+1)+O(1)$; [6] improved it to $I+\log(I+1)+3.870$; [7] improved it to $I+\log(I+1)+3.732$; and [8] improved it to $I+\log(I+2)+2$. See [11] for a review. In Section VI, we show a new bound $I+\log(I+5.51)+1.06$, which improves upon [7] for every $I$, and improves upon [8] for $I\geq 2$. These improved bounds show the usefulness of $\Lambda(X)$ as a technical tool.
Notions that can be interpreted as difference have been studied systematically in [12]. It was argued that apart from the strong functional representation lemma, the Körner graph entropy [13] and the maximum rate for perfect privacy [14] can be treated as the difference between $X$ and $Y$. The minimum entropy coupling [15, 16, 17, 18] has also been related to the difference between information.
The set corresponding to the cell “$X\backslash Y$” in the I-measure [19] has been studied in [20, 21], though this “$X\backslash Y$” is not actually a random variable.
Notations
Throughout this paper, all random variables are assumed to be discrete (with finite or countable support) unless otherwise stated. Entropy is in bits, and $\log$ is to the base $2$. Write :=$\mathbb{N}:=\{1,2,\ldots\}$, $[n]:=\{1,\ldots,n\}$ (let $[\infty]:=\mathbb{N}$). For a random variable $X$, write $\mathcal{X}$ for the set it lies in, and $p_{X}$ for its probability mass function (pmf). For random variables $X,Y$, we say that they are (informationally) equivalent, denoted as $X\stackrel{{\scriptstyle\iota}}{{=}}Y$, if $H(X|Y)=H(Y|X)=0$. Denote the constant random variable as $\emptyset$ (so $X\stackrel{{\scriptstyle\iota}}{{=}}\emptyset$ means that $X$ is a constant). Write $X^{n}=(X_{1},\ldots,X_{n})$. The min-entropy is defined as $H_{\infty}(X):=-\log\max_{x}p_{X}(x)$.
For a pmf $p$ over $\mathcal{X}$, write $p^{\downarrow}$ for a pmf over $\mathbb{N}$, where $p^{\downarrow}(i)$ is the $i$-th entry of $(p(x))_{x\in\mathcal{X}}$ when sorted in descending order ($p^{\downarrow}(i)=0$ if $i>|\mathcal{X}|$). Given pmfs $p,q$, we say that $p$ majorizes $q$, written as $p\succeq q$, if $\sum_{i=1}^{k}p^{\downarrow}(i)\geq\sum_{i=1}^{k}q^{\downarrow}(i)$ for every $k\in\mathbb{N}$ [22]. A function $\Lambda$ mapping pmfs to real numbers is Schur concave if $p\succeq q$ implies $\Lambda(p)\leq\Lambda(q)$. For example, Shannon entropy is Schur concave.
II Discrete Layered Entropy
We now define the central quantity of this paper.
Definition 1 (Discrete layered entropy).
The discrete layered entropy of a probability mass function $p$ is defined as
$\Lambda(p):=\sum_{i=1}^{\infty}p^{\downarrow}(i)\left(i\log i-(i-1)\log(i-1)\right).$ (3)
Recall that $p^{\downarrow}(i)$ is the $i$-th entry of $(p(x))_{x\in\mathcal{X}}$ when sorted in descending order, and we assume $00\log 0=$ 0. We write $\Lambda(X):=\Lambda(p_{X})$. The conditional discrete layered entropy of a random variable $X$ given another random variable $Y$ is
$\Lambda(X|Y):=\mathbb{E}_{Y}\left[\Lambda(p_{X|Y}(\cdot|Y))\right].$
We use the name “discrete layered entropy” since $\Lambda(p)$ is the discrete analogue of the (continuous) layered entropy studied in [1, 2], which is shown in (5). [^2] The notation “$\Lambda(p)$” is chosen for two reasons—both “layered” and “lambda” starts with “la”, and the shape “$\Lambda$” mimics the shape of the function $\Lambda(p)$ (a piecewise-linear concave function) when $p$ is binary. We give several alternative definitions for $\Lambda(p)$. The proof is given in Appendix -A.
Proposition 2 (Alternative definitions).
We have
  1. 1.
    (Integral form)
    $\Lambda(p)=\int_{0}^{\infty}p^{\downarrow}(\lceil t\rceil)\log(et)\mathrm{d}t.$ (4)
  2. 2.
    (Layered form)
    $\Lambda(p)=\int_{0}^{1}|\{x:\,p(x)>t\}|\cdot\log|\{x:\,p(x)>t\}|\mathrm{d}t.$ (5)
  3. 3.
    (Concave envelope of min-entropy)
    $\Lambda(X)=\max_{p_{Y|X}}H_{\infty}(X|Y),$ (6)
    where $H_{\infty}(X|Y):=\mathbb{E}_{Y}[-\log\max_{x}p_{X|Y}(x|Y)]$ is the conditional min-entropy. In other words, $\Lambda$ is the upper concave envelope of $H_{\infty}$. [^3] This means $\Lambda(p)=\mathrm{min}_{\tilde{\Lambda}}\tilde{\Lambda}(p)$ where the minimum is over concave functions $\tilde{\Lambda}$ satisfying $\tilde{\Lambda}(p)\geq H_{\infty}(p)$ for all $p$.
  4. 4.
    (Concave envelope of log cardinality)
    $\Lambda(X)=\max H(X|Y),$ (7)
[…]
  5. 5.
    (Linear programming form)
    $\Lambda(X)=\max\mathbb{E}[\log K],$ (8)
    where the maximum is over joint probability mass functions $p_{X,K}$ over $\mathcal{X}\times[|\mathcal{X}|]$ with an $X$-marginal that coincides with $p_{X}$, and satisfying that $p_{X,K}(x,k)\leq p_{K}(k)/k$ for all $x,k$, where $p_{K}(k)$ is the $K$-marginal of $p_{X,K}$. This means $\Lambda(X)$ can be formulated as a maximization in a linear program when $\mathcal{X}$ is finite.
Another alternative definition of $\Lambda(X)$ will be given in Theorem 8. We then show some basic properties of $\Lambda(X)$. The proof is given in Appendix -B.
Proposition 3 (Basic properties).
We have, for every random variables $X\in\mathcal{X}$, $Y\in\mathcal{Y}$,
  1. 1.
    $H_{\infty}(X)\leq\Lambda(X)\leq H(X)$. For each inequality, equality holds if and only if $X$ is uniformly distributed.
  2. 2.
    (Concavity) $\Lambda(p)$ is a concave function over probability mass functions $p$. Equivalently, $\Lambda(X|Y)\leq\Lambda(X)$.
  3. 3.
    (Schur concavity) $\Lambda(p)$ is Schur concave, and hence $\Lambda(Y)\leq\Lambda(X)$ if $H(Y|X)=0$.
  4. 4.
    (Monotone linearity) $\Lambda(p)$ is a linear function over the convex space of nondecreasing probability mass functions $p:\mathbb{N}\to\mathbb{R}$. Equivalently, if $X\in\mathbb{N}$, and for every fixed $y$, $x\mapsto p_{X|Y}(x|y)$ is a nondecreasing function, then $\Lambda(X|Y)=\Lambda(X)$. (Overall, $\Lambda(p)$ is a piecewise linear function over the space of not-necessarily-nondecreasing probability mass functions.)
  5. 5.
    (Superadditivity) If $X,Y$ are independent, then
    $\Lambda(X,Y)\geq\Lambda(X)+\Lambda(Y).$
    Equality holds if and only if at least one of $X,Y$ is uniformly distributed.
  6. 6.
    (Bounded increase) For any $X,Y$,
    $\Lambda(X,Y)\leq\Lambda(X)+\log|\mathcal{Y}|.$
    Equality holds if and only if $Y$ is uniformly distributed and is independent of $X$.
One of the most useful property of $\Lambda$ is that $\Lambda(X)$ is close to $H(X)$ within a logarithmic gap. The proof is in Appendix -C.
Proposition 4 ($\Lambda(X)\approx H(X)$).
For every discrete $X$,
$\Lambda(X)\leq H(X)\leq\Lambda(X)+\log\left(1+\frac{\Lambda(X)}{e\eta}\right)+\eta$ (9)
for every $\eta>0$. In particular, taking $\eta=\log e$ gives a bound good for large $\Lambda(X)$:
$\Lambda(X)\leq H(X)\leq\Lambda(X)+\log\left(1+\frac{\Lambda(X)}{e\log e}\right)+\log e.$
Taking $\eta=\sqrt{\frac{\Lambda(X)\log e}{e}}$ gives a bound good for small $\Lambda(X)$:
$\Lambda(X)\leq H(X)$ $\leq\Lambda(X)+2\sqrt{\Lambda(X)e^{-1}\log e}.$
Note that $\log(1+\Lambda/(e\eta))+\eta$ (where $\Lambda=\Lambda(X)$) is minimized at $\eta=\frac{1}{2e}(\sqrt{\Lambda^{2}+4e\Lambda\log e}-\Lambda)$. This gives the best bound for (9), but is a little unwieldy.
Proposition 4 shows that $\Lambda$ is a good approximate of $H$, and can be used in place of $H$ when the properties of $\Lambda$ are more desirable. For example, consider the entropy maximization problem $\max_{p\in\mathcal{S}}H(p)$ where $\mathcal{S}$ is a convex polytope of probability vectors. This cannot be solved via linear programming since $H(p)$ is not piecewise linear/affine. Moreover, $H(p)$ has undefined gradient when there is a zero entry in $p$, making it difficult to solve the problem via gradient-based methods. On the other hand, the layered entropy maximization problem $\max_{p\in\mathcal{S}}\Lambda(p)$ can be solved via linear programming by (8). By Proposition 4, the optimal values of the two problems are close within a logarithmic gap, making $\Lambda$ a good substitute for $H$. In the following sections, we will see more theoretical properties of $\Lambda$ that are not satisfied by $H$.
III Relation to One-to-one Non-prefix-free Codes
A one-to-one (non-prefix-free) code is an injective function $f:\mathcal{X}\to\{0,1\}^{*}$ which is not subject to the prefix-free requirement [3, 4, 5]. The optimal expected length of a one-to-one encoding of $X$ is [3, 4]
$L(X):=\sum_{i=1}^{\infty}p_{X}^{\downarrow}(i)\lfloor\log i\rfloor.$
We show that $L(X)$ is approximated by $\Lambda(X)$ within $2$ bits. This is in contrast to $H(X)$ which approximates the expected length of prefix-free codes. The proof is in Appendix -D.
Proposition 5.
We have
$\Lambda(X)-2<L(X)\leq\Lambda(X).$
Using Propositions 4 and 5, we can show that the optimal expected length of one-to-one codes is at least $H(X)-\log(H(X)+1)-O(1)$, giving a similar (and slightly weaker) bound compared to [3, 4]. By [5], when $X^{n}=(X_{1},\ldots,X_{n})$ is an i.i.d. sequence following $p_{X}$, as $n\to\infty$,
$\Lambda(X^{n})=\begin{cases}nH(X)&\text{if}\;X\;\text{is uniform},\\ nH(X)-\frac{\log n}{2}+O(1)&\text{if}\;\text{nonuniform}.\end{cases}$
One may raise the question—why should we study $\Lambda$ instead of $L$ which is exactly the optimal length? Indeed, $L$ also satisfies some properties of $\Lambda$, such as concavity, monotone linearity and the conditional property to be discussed in Proposition 9. One reason for preferring $\Lambda$ is that $\Lambda(X)$ approximates $H(X)$ better than $L(X)$. We have $L(X)\leq\Lambda(X)\leq H(X)$, and $\Lambda(X)=H(X)$ whenever $X$ is uniform, whereas the expression of $L(X)$ for uniform $X$ is complicated. This not only allows $\Lambda$ to satisfy more elegant theoretical properties compared to $L$ (Theorems 8, 11, 12, and some properties in Proposition 3), but also makes $\Lambda$ a more useful tool for approximation tasks (e.g., Theorem 16).
IV Conditional Compression and “Three Conditional Entropies”
IV-A Conditional Compression
Given random variables $X,Y$, we are interested in finding “the information in $(X,Y)$ that is not in $Y$”. Intuitively, it is the smallest $U$ such that $X$ can be recovered using $(Y,U)$. It is formally defined below.
Definition 6 (Conditional compression).
[…]
The distribution of a conditional compression can be found.
Proposition 7.
The probability mass function of any conditional compression $U$ of $X$ given $Y$ must satisfy
$p_{U}^{\downarrow}(i)=\mathbb{E}_{Y}\big{[}p_{X|Y}^{\downarrow}(i|Y)\big{]},$
for $i\in\mathbb{N}$, where $p_{X|Y}^{\downarrow}(i|y)$ is the $i$-th entry of $(p_{X|Y}(x|y))_{x\in\mathcal{X}}$ when sorted in descending order.
Proof:
We can find a conditional compression in the following way. For each $y$, we sort the values of $(p_{X|Y}(x|y))_{x\in\mathcal{X}}$ in descending order to obtain $p_{X|Y}(x_{y}^{(1)}|y)\geq p_{X|Y}(x_{y}^{(2)}|y)\geq\cdots$, where $x_{y}^{(1)},x_{y}^{(2)},\ldots\in\mathcal{X}$ are distinct values. Take $U\in\mathbb{N}$ to be the value that satisfies $X=x_{Y}^{(U)}$, i.e., $U$ is the ranking of $p_{X|Y}(X|Y)$ among $(p_{X|Y}(x|Y))_{x\in\mathcal{X}}$ ($p_{X|Y}(X|Y)$ is the $U$-th largest among $p_{X|Y}(x|Y)$ for $x\in\mathcal{X}$). We then have $p_{U}(i)=\mathbb{E}_{Y}[p_{X|Y}(x_{Y}^{(i)}|Y)]=\mathbb{E}_{Y}[p_{X|Y}^{\downarrow}(i|Y)]$.
To show that this is indeed a conditional compression (it minimizes $H(U)$), consider any other $V$ satisfying $H(X|Y,V)=0$. Hence, $p_{X|Y}(\cdot|y)\succeq p_{V|Y}(\cdot|y)$ for every $y$, i.e., $\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|y)\geq\sum_{i=1}^{k}p_{V|Y}^{\downarrow}(i|y)$. This implies
$\sum_{i=1}^{k}p_{V}^{\downarrow}(i)$ $\leq\mathbb{E}\Big{[}\sum_{i=1}^{k}p_{V|Y}^{\downarrow}(i|Y)\Big{]}$
$\leq\mathbb{E}\Big{[}\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|Y)\Big{]}=\sum_{i=1}^{k}p_{U}(i),$
and $p_{U}\succeq p_{V}$, and hence $H(U)\leq H(V)$. Equality holds if and only if $p_{V}^{\downarrow}(i)=p_{U}(i)$, i.e., $V$ has the same distribution as $U$ up to relabeling. Therefore, any conditional compression has the same distribution $p_{U}$ up to relabeling. ∎
The reason for the notation $X\backslash Y$ is that the conditional compression shares a number of properties with set difference. Firstly, $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}\emptyset$ (i.e., $X\backslash Y$ is a constant) if $H(X|Y)=0$, corresponding to the fact that $A\backslash B=\emptyset$ for sets $A,B$ if $A\subseteq B$. Secondly, $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$ if $X\perp\!\!\!\perp Y$, analogous to the fact that $A\backslash B=A$ if $A\cap B=\emptyset$. This property is the reason of the tie-breaking rule (minimizing $H(X|U)$ in case of a tie of minimizing $H(U)$) in the definition of $X\backslash Y$ [^6] Although there may be multiple conditional compressions in case there are ties among $(p_{X|Y}(x|Y))_{x\in\mathcal{X}}$, we are usually only interested in the marginal distribution $p_{U}$ so it does not matter which one we consider (they all have the same distribution). Nevertheless, in order to justify the notation $X\backslash Y$, we select a “canonical” conditional compression $X\backslash Y$ in Definition 6 to be the one that minimizes $H(X|U)$, to ensure that $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$ if $X\perp\!\!\!\perp Y$.
The converse of the first property holds ($X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}\emptyset$ if and only if $H(X|Y)=0$), but the converse of the second property is false, i.e., $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$ does not imply $I(X;Y)=0$. In fact, $H(X)-\Lambda(X)$ is the maximum violation, i.e., $H(X)-\Lambda(X)$ is the largest possible $I(X;Y)$ subject to $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$. This gives a simple alternative definition of $\Lambda(X)$ given in Theorem 8. By Proposition 4, if $H(X\backslash Y)=H(X)$, then $I(X;Y)\leq H(X)-\Lambda(X)\leq O(\log H(X))$, so the converse does approximately hold. The proof of Theorem 8 is in Appendix -E. In Section IV-B, we will discuss the operational meaning of this definition.
Theorem 8 (Alternative definition of $\Lambda(X)$).
We have
$\Lambda(X)$ $=\!\min_{p_{Y|X}:\,H(X\backslash Y)=H(X)}\!\!H(X|Y)=\!\min_{p_{Y|X}:\,X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X}\!\!\!H(X|Y).$
Also, if $H(X\backslash Y)=H(X)$, then $\Lambda(X\backslash Y)=\Lambda(X)$.
An elegant property of $\Lambda(X)$ is that $\Lambda(X|Y)=\mathbb{E}_{Y}[\Lambda(p_{X|Y}(\cdot|Y)]$ coincides with $\Lambda(X\backslash Y)$. Therefore, $\Lambda(X|Y)$ is indeed the discrete layered entropy of “the random variable $X|Y$” which is formally given as $X\backslash Y$.
Proposition 9 (Conditioning property).
$\Lambda(X|Y)=\Lambda(X\backslash Y).$
Equivalently,
$\Lambda(X|Y)=\min_{p_{U|X,Y}:\,H(X|Y,U)=0}\Lambda(U).$ (10)
We will see in Section IV-C that the conditioning property is actually the defining property of the discrete layered entropy, in the sense that if the function $\Lambda$ satisfies the conditioning property and $\Lambda(X)=\log|\mathcal{X}|$ if $X$ is uniformly distributed, then $\Lambda$ must be the discrete layered entropy.
We now have three different notions of “conditional entropy”, namely $H(X|Y)$, $H(X\backslash Y)$ and $\Lambda(X|Y)=\Lambda(X\backslash Y)$. Although Shannon entropy does not satisfy the conditioning property, i.e., $H(X\backslash Y)$ is generally not equal to $H(X|Y)$, we can show that $H(X\backslash Y)\approx H(X|Y)$ via the discrete layered entropy. By Propositions 3 and 4,
$\Lambda(X|Y)$ $\leq H(X|Y)\leq H(X\backslash Y)$
$\;\;\leq\Lambda(X|Y)+\log\left(1+\frac{\Lambda(X|Y)}{e\eta}\right)+\eta+1.$ (11)
Therefore, the three “conditional entropies” are close within a logarithmic gap. We highlight the following theorem that follows directly from (11).
Theorem 10.
For every $\eta>0$,
$H(X|Y)\leq H(X\backslash Y)\leq H(X|Y)+\log\left(1+\frac{H(X|Y)}{e\eta}\right)+\eta.$
Theorem 10 can also be stated as: for every $X,Y$, there exists $U$ such that $H(X|Y,U)=H(U|X,Y)=0$, and for every $\eta>0$, $H(U)\leq H(X|Y)+\log(1+H(X|Y)/(e\eta))+\eta$. This is an example of a nonlinear existential information inequality [23, 24, 25]. It is perhaps interesting that such a simple (and novel, to the best of the author’s knowledge) fact about Shannon entropy can be proved via the properties of discrete layered entropy $\Lambda(X)$. It shows the usefulness of $\Lambda(X)$ as a tool for proving results about Shannon entropy.
Also, unlike set difference where $A\backslash B$ and $B$ are disjoint, we do not have $I(X\backslash Y;Y)=0$. Nevertheless, Theorem 10 implies that this property approximately holds, in the sense that $I(X\backslash Y;Y)\leq\log(1+H(X|Y)/(e\eta))+\eta$ is small. [^7] Compare $X\backslash Y$ to the strong functional representation lemma (SFRL) [6, 8]: for every $X,Y$, there exists $S$ such that $I(S;Y)=H(X|Y,S)=0$ and $H(X|S)\leq I(X;Y)+\log(I(X;Y)+2)+2$. SFRL guarantees $I(S;Y)=0$, whereas we only have $I(X\backslash Y;Y)\approx 0$. Nevertheless, SFRL does not guarantee $H(S)\approx H(X|Y)$ (it only has $I(S;X,Y)=H(X|Y)$; the construction in [6] can have extremely large $H(S)$), whereas we have $H(X\backslash Y)\approx H(X|Y)$.
In the next subsection, we will give operational meanings to the three “conditional entropies”, and see how (11) and Theorem 10 are relevant.
IV-B Conditional Variable-Length Encoding
Consider a one-shot variable-length encoding setting, where there is a source symbol $X$ that is correlated with a side information $Y$. The encoder observes $X,Y$, and sends a variable-length description $M\in\{0,1\}^{*}$ to the decoder. The decoder observes $Y,M$, and has to recover $X$ losslessly, i.e., $H(X|Y,M)=0$. The goal is to make the expected length $\mathbb{E}[|M|]$ as small as possible.
If there is no prefix-free requirement on $M$, the task is straightforward—given that $Y=y$, the encoder encodes $X$ into $M$ using the best non-prefix-free one-to-one code designed for the distribution $p_{X|Y}(\cdot|y)$. Let $\ell_{\mathrm{n}}^{*}$ be the smallest possible $\mathbb{E}[|M|]$ in this case. By Proposition 5, $\Lambda(X|Y)-2<\ell_{\mathrm{n}}^{*}\leq\Lambda(X|Y)$. As usual for a non-prefix-free codes, the decoder cannot synchronize with the encoder, i.e., the decoder does not know where $M$ ends, unless external help is available (e.g., when $|M|$ is already given by the metadata of the communication protocol).
For the sake of synchronization, given that $Y=y$, this task is conventionally performed by encoding $X$ conditional on $Y=y$ into $M$, using a prefix-free code designed for the distribution $p_{X|Y}(\cdot|y)$. This gives a “conditionally prefix-free code”, i.e., $M\in\mathcal{C}_{Y}$, where $\mathcal{C}_{y}\subseteq\{0,1\}^{*}$ is a prefix-free codebook for every $y\in Y$. Let $\ell_{\mathrm{c}}^{*}$ be the smallest possible $\mathbb{E}[|M|]$ in this case. Using Huffman coding [26], $H(X|Y)\leq\ell_{\mathrm{c}}^{*}<H(X|Y)+1$.
A shortcoming of this approach is that this “conditional prefix-free code” is not a prefix-free code to a party that does not know $Y$. This party only knows $M\in\bigcup_{y\in\mathcal{Y}}\mathcal{C}_{y}$ which may not be a prefix-free codebook. For example, consider $X=\{1,2,3\}$, $Y=\{1,2\}$, $p_{X|Y}(1|1)=p_{X|Y}(2|1)=p_{X|Y}(1|2)=1/2$, $p_{X|Y}(2|2)=p_{X|Y}(3|2)=1/4$. The optimal conditional prefix-free code $M=f(X,Y)$ would be $f(1,1)=0$, $f(2,1)=1$, $f(1,2)=0$, $f(2,2)=10$, $f(3,2)=11$, which is not prefix-free. This shortcoming can be significant if the decoder does not always know $Y$. Without knowing $Y$, the decoder not only cannot decode $X$, but also cannot know where the bit sequence ends, and becomes desynchronized with the encoder.
For example, consider a video streaming application where the encoder sends the frames of a video to the decoder. Since consecutive frames are correlated, it is reasonable for the encoder to encode each frame conditional on the previous frame to reduce the encoding size. However, if there is a packet loss, the decoder will be unable to decode all subsequent frames that are encoded conditionally on the previous frames. Therefore, the encoder would occasionally encode some frames (the keyframes) by themselves without conditioning on the previous frames, so the decoder can decode these keyframes even if the previous frames are lost.
Now, consider the scenario where there is a packet loss and the decoder cannot decode the current frame. A new packet arrives which contains some conditionally-encoded frames, with one keyframe in the middle of the packet. If a conditional prefix-free code is used to encode the conditionally-encoded frames, then the decoder will become desynchronized and fail to decode the keyframe. In comparison, if an unconditional prefix-free code is used (i.e., $M$ lies in a prefix-free codebook that does not depend on $Y$), then the decoder will be unable to decode the conditionally-encoded frames, but still be able to keep synchronous when it scans through the packet, and hence can decode the keyframe.
Hence, it is of interest to study the setting where $M$ is in an unconditionally prefix-free codebook, where we let $\ell_{\mathrm{u}}^{*}$ be the smallest possible $\mathbb{E}[|M|]$. Using Huffman coding [26], $\ell_{\mathrm{u}}^{*}$ is within $1$ bit from $\min_{U:\,H(X|Y,U)=0}H(U)=H(X\backslash Y)$. We summarize the optimal expected lengths in the three settings:
  • For non-prefix-free one-to-one, $\ell_{\mathrm{n}}^{*}\approx\Lambda(X|Y)$:
    $\Lambda(X|Y)-2<\ell_{\mathrm{n}}^{*}\leq\Lambda(X|Y).$ (12)
  • For conditionally prefix-free, $\ell_{\mathrm{c}}^{*}\approx H(X|Y)$:
    $H(X|Y)\leq\ell_{\mathrm{c}}^{*}<H(X|Y)+1.$ (13)
  • For unconditionally prefix-free, $\ell_{\mathrm{u}}^{*}\approx H(X\backslash Y)$:
    $H(X\backslash Y)\leq\ell_{\mathrm{u}}^{*}<H(X\backslash Y)+1.$ (14)
The three “conditional entropies” $\Lambda(X|Y)\leq H(X|Y)\leq H(X\backslash Y)$ are close within a logarithmic gap, as shown in (11).
To study the difference between conditionally (13) and unconditionally (14) prefix-free encoding, it is of interest to study the region of possible values of the pair $(H(X|Y),H(X\backslash Y))\in\mathbb{R}^{2}$. Given $X$, define
$\mathcal{R}(X):=\bigcup_{p_{Y|X}}\left\{(H(X|Y),H(X\backslash Y))\right\}.$
An inner bound of $\mathcal{R}(X)$ is given by the diagonal line, that is, $(t,t)\in\mathcal{R}(X)$ for every $0\leq t\leq H(X)$. [^8] To show this, consider a distribution $q$ that majorizes $p_{X}$ [22] with $H(q)=t$. We can construct random variables $U,Y$ such that $U\sim q$, $U$ is independent of $Y$, and $H(X|U,Y)=H(U|X,Y)=0$. This gives $H(X|Y)=H(U|Y)=H(U)=t$. Since $p_{X|Y}(\cdot|y)$ contains the same entries (possibly reordered) as $q(\cdot)$, $U$ is a conditional compression of $X$ given $Y$, and hence $H(X\backslash Y)=H(U)=t$. Theorem 10 gives an outer bound of $\mathcal{R}(X)$, showing that $\mathcal{R}(X)$ cannot be too far from the diagonal line.
Figure 2: Illustration of $\mathcal{R}(X)=\bigcup_{p_{Y|X}}\{(H(X|Y),H(X\backslash Y))\}$ showing the extreme points $(\Lambda(X),H(X))$ and $(H(X),H(X))$.
An interesting extreme point of $\mathcal{R}(X)$ is the minimum of $H(X|Y)$ subject to the contraint that $H(X\backslash Y)$ is maximized ($H(X\backslash Y)=H(X)$). The $Y$ that attains this minimum is the most “conditionally useful” (i.e., giving the shortest conditionally prefix-free encoding length $H(X|Y)$) among “unconditionally useless” side information (i.e., the unconditionally prefix-free encoding length $H(X\backslash Y)$ is the same as if $Y$ is absent). By Theorem 8, this minimum is given by $\Lambda(X)$. Hence, we have another operational meaning of $\Lambda(X)$: “if we are to design a side information $Y$ that provides no benefit to a party that encodes $X$ using an unconditionally prefix-free code, how useful can it be to a party that uses a conditionally prefix-free code?”
Even more strangely, $\Lambda(X)$ is also the $\Lambda(X|Y)$ (non-prefix-free length (12)) when $H(X|Y)$ is maximized, or when $H(X\backslash Y)$ is maximized (by Theorem 8). So,
$\Lambda(X)$ $=\min_{p_{Y|X}:\,H(X\backslash Y)=H(X)}H(X|Y)$
$=\min_{p_{Y|X}:\,H(X\backslash Y)=H(X)}\Lambda(X|Y)$
$=\min_{p_{Y|X}:\,H(X|Y)=H(X)}\Lambda(X|Y),$
and hence every pair of the three settings (12), (13), (14) have the same gap in this sense. An intuitive explanation for this curious coincidence is left for future studies.
IV-C Characterizations of $\Lambda(X)$ via the Conditioning Property
Proposition 8 gives a definition of $\Lambda(X)$ using only Shannon entropy. In this subsection, we study some more axiomatic definitions of $\Lambda(X)$ via the conditioning property $\Lambda(X|Y)=\Lambda(X\backslash Y)$ (Proposition 9). We first show that any function $\tilde{\Lambda}$ satisfying the conditioning property and $\tilde{\Lambda}(X)=\log|\mathcal{X}|$ for uniform $X$ must be the discrete layered entropy. The proof is given in Appendix -F.
Theorem 11.
If $\tilde{\Lambda}$ is a function mapping discrete distributions to nonnegative real numbers that satisfies the conditioning property (i.e., $\tilde{\Lambda}(X\backslash Y)=\tilde{\Lambda}(X|Y):=\mathbb{E}_{Y}[\tilde{\Lambda}(p_{X|Y}(\cdot|Y))]$), $\tilde{\Lambda}(X)=\log|\mathcal{X}|$ whenever $X$ is uniformly distributed over $\mathcal{X}$, and $\tilde{\Lambda}(X)=\tilde{\Lambda}(Y)$ whenever $X\stackrel{{\scriptstyle\iota}}{{=}}Y$ (i.e., $\tilde{\Lambda}$ is invariant under relabeling), then $\tilde{\Lambda}(X)=\Lambda(X)$.
Compare this with the axiomatic characterization of Shannon entropy in [27]: if $\tilde{H}$ satisfy the subadditivity property ($\tilde{H}(X,Y)\geq\tilde{H}(X)+\tilde{H}(Y)$), additivity property ($\tilde{H}(X,Y)=\tilde{H}(X)+\tilde{H}(Y)$ if $X\perp\!\!\!\perp Y$), continuity with respect to the probability mass function, $\tilde{H}(X)=1$ if $X\sim\mathrm{Bern}(1/2)$, and $\tilde{H}(X)=\tilde{H}(Y)$ whenever $X\stackrel{{\scriptstyle\iota}}{{=}}Y$, then $\tilde{H}$ is the Shannon entropy $H$. We can see that additivity, subadditivity and continuity are the defining properties of Shannon entropy, whereas conditioning is the defining property of discrete layered entropy.
We then study another characterization of $\Lambda(X)$. Recall that $H(X)$ does not satisfy the conditioning property. Nevertheless, we can ask what is the best under-approximation of $H(X)$ that satisfy the conditioning property. The answer is given by $\Lambda(X)$. The proof is in Appendix -G.
Theorem 12.
If $\tilde{\Lambda}$ is a function mapping discrete distributions to nonnegative real numbers that satisfies the conditioning property, $\tilde{\Lambda}(X)\leq H(X)$ for every $X$, [^9] Actually, we only need $\tilde{\Lambda}(X)\leq\log|\mathcal{X}|$ whenever $X$ is uniform. and $\tilde{\Lambda}(X)=\tilde{\Lambda}(Y)$ whenever $X\stackrel{{\scriptstyle\iota}}{{=}}Y$, then $\tilde{\Lambda}(X)\leq\Lambda(X)$ for every $X$. Equivalently, $\Lambda(X)$ admits the following characterization:
$\Lambda(X)=\mathrm{max}_{\tilde{\Lambda}}\tilde{\Lambda}(X),$
where the maximum is over functions $\tilde{\Lambda}$ that satisfies the conditioning property, $\tilde{\Lambda}(Y)\leq H(Y)$ for every $Y$, and $\tilde{\Lambda}(Y)=\tilde{\Lambda}(Z)$ whenever $Y\stackrel{{\scriptstyle\iota}}{{=}}Z$.
V Discrete Rényi Layered Entropy
Rényi entropy [28] is a generalization of Shannon entropy, defined as (where $\alpha\in(0,\infty)\backslash\{1\}$ is the order)
$H_{\alpha}(p):=\frac{1}{1-\alpha}\log\sum_{x}(p(x))^{\alpha}.$
Also, define $H_{\alpha}(X):=\lim_{\alpha^{\prime}\to\alpha}H_{\alpha^{\prime}}(X)$ for $\alpha\in\{0,1,\infty\}$, which gives $H_{1}(p)=H(p)$, $H_{0}(p)=\log|\{x:\,p(x)>0\}|$, and $H_{\infty}(p)=-\log\max_{x}p(x)$. In this section, we generalize the discrete layered entropy to the discrete Rényi layered entropy in an analogous manner.
Definition 13 (Discrete Rényi layered entropy).
Given $\alpha\in(0,\infty)\backslash\{1\}$, the discrete Rényi layered entropy of order $\alpha$ of a probability mass function $p$ is defined as
$\Lambda_{\alpha}(p):=\frac{1}{1/\alpha-1}\log\sum_{i=1}^{\infty}p^{\downarrow}(i)\big{(}i^{1/\alpha}-(i-1)^{1/\alpha}\big{)},$ (15)
where $p^{\downarrow}(i)$ is the $i$-th entry of $(p(x))_{x\in\mathcal{X}}$ when sorted in descending order. Also, define $\Lambda_{\alpha}(X):=\lim_{\alpha^{\prime}\to\alpha}\Lambda_{\alpha^{\prime}}(X)$ for $\alpha\in\{0,1,\infty\}$, which can be evaluated as $\Lambda_{0}(p)=H_{0}(p)$, $\Lambda_{\infty}(p)=H_{\infty}(p)$, and $\Lambda_{1}(p)=\Lambda(p)$.
We present some properties of $\Lambda_{\alpha}(p)$. The proof is in Appendix -H.
Proposition 14.
We have the following for $\alpha\in[0,\infty]$:
  • $\Lambda_{\alpha}(X)$ is non-increasing in $\alpha$, and $\Lambda_{\alpha}(X)\leq H_{\alpha}(X)$.
  • When $X$ is uniformly distributed, $\Lambda_{\alpha}(X)=H_{\alpha}(X)=\log|\mathcal{X}|$.
  • If $X\in\mathbb{N}$, then
    $\Lambda_{\alpha}(X)\leq\frac{1}{1/\alpha-1}\log\mathbb{E}\left[X^{1/\alpha}-(X-1)^{1/\alpha}\right].$
    In particular, $\Lambda_{1/2}(X)\leq\log(2\mathbb{E}[X]-1)$.
Remark 15.
Given (6), another way to modify Rényi entropy to incorporate $\Lambda(X)$ is to study the concave envelope of the Rényi entropy
$\overline{H}_{\alpha}(X):=\max_{p_{Y|X}}H_{\alpha}(X|Y),$
where $H_{\alpha}(X|Y):=\mathbb{E}_{Y}[H_{\alpha}(p_{X|Y}(\cdot|Y))]$. Since $H_{\alpha}$ is already concave for $\alpha\leq 1$, we have $\overline{H}_{\alpha}(X)=H_{\alpha}(X)$ for $\alpha\leq 1$. We also have $\overline{H}_{\infty}(X)=\Lambda(X)$. Therefore, we can use $\overline{H}_{\alpha}(X)$ to interpolate between $H(X)$ and $\Lambda(X)$. The properties of $\overline{H}_{\alpha}$ are left for future studies.
VI Strong Functional Representation Lemma
Using the properties of $\Lambda$, we can give a tighter bound for the strong functional representation lemma which improves upon [6, 7, 8] for $I(X;Y)\geq 2$. If we optimize over $\eta$, then this improves upon [8] for $I(X;Y)\geq 0.7$. The proof is in Appendix -I.
Theorem 16.
For every (not necessarily discrete) random variables $X,Y$, there exists a (not necessarily discrete) random variable $S$ such that $S$ is independent of $X$, $H(Y|X,S)=0$, and
$\Lambda(Y|S)\leq I(X;Y)+\log 3.$
Hence, for every $\eta>0$,
$H(Y|S)\leq I(X;Y)+\log\left(I(X;Y)+\log 3+e\eta\right)+\log\frac{3}{e\eta}+\eta.$ (16)
In particular, substituting $\eta=\log e$ gives
$H(Y|S)\leq I(X;Y)+\log\left(I(X;Y)+5.51\right)+1.06.$ (17)
Refer to Figure 3 for a comparison between Theorem 16 and previous bounds. It includes the plot of the bound $I+\log(I+1)+3.732$ (writing $I=I(X;Y)$) in [7]; $I+\log(I+2)+2$ in [8]; our new bound (17); and our new bound (16) when optimized over $\eta$ (the optimal $\eta$ is $\frac{1}{2e}(\sqrt{\Lambda^{2}+4e\Lambda\log e}-\Lambda)$ where $\Lambda=I+\log 3$).
Figure 3: Comparison between Theorem 16 and previous bounds [7, 8] for the strong functional representation lemma.
VII Acknowledgement
This work was partially supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No.s: CUHK 24205621 (ECS), CUHK 14209823 (GRF)].
-A Proof of Proposition 2
(4) follows from $i\log i-(i-1)\log(i-1)=\int_{i-1}^{i}\log(et)\mathrm{d}t$. For (5),
$\Lambda(p)$ $=\sum_{i=1}^{\infty}p^{\downarrow}(i)\left(i\log i-(i-1)\log(i-1)\right)$
$=\sum_{i=1}^{\infty}\int_{0}^{1}\mathbf{1}\{p^{\downarrow}(i)>t\}\left(i\log i-(i-1)\log(i-1)\right)\mathrm{d}t$
$=\int_{0}^{1}\sum_{i=1}^{\infty}\mathbf{1}\{p^{\downarrow}(i)>t\}\left(i\log i-(i-1)\log(i-1)\right)\mathrm{d}t$
$=\int_{0}^{1}|\{x:\,p(x)>t\}|\cdot\log|\{x:\,p(x)>t\}|\mathrm{d}t.$
We then prove
$\Lambda(p)=\mathrm{min}_{\tilde{\Lambda}}\tilde{\Lambda}(p),$ (18)
where the minimum is over all concave functions $\tilde{\Lambda}$ satisfying $\tilde{\Lambda}(X)\geq\log|\mathcal{X}|$ when $X$ is uniformly distributed. Assume $X\in\mathbb{N}$ and $p_{X}(x)$ is in descending order without loss of generality. Note that
$p_{X}(x)=\sum_{i=1}^{\infty}i(p_{X}(i)-p_{X}(i+1))\cdot\frac{\mathbf{1}\{x\leq i\}}{i}$
is a convex combination of uniform distributions $\mathrm{Unif}([i])$. Hence, for every concave $\tilde{\Lambda}$ satisfying that $\tilde{\Lambda}(\mathrm{Unif}([i]))\geq\log i$, we have
$\tilde{\Lambda}(X)$ $\geq\sum_{i=1}^{\infty}i(p_{X}(i)-p_{X}(i+1))\cdot\log i$
$=\Lambda(X).$
To show that there exists $\tilde{\Lambda}$ satisfying $\tilde{\Lambda}(Y)\geq\log|\mathcal{Y}|$ whenever $Y$ is uniform and $\tilde{\Lambda}(X)=\Lambda(X)$ for this fixed $p_{X}$, we take
$\tilde{\Lambda}(p)=\sum_{i=1}^{\infty}p(i)\left(i\log i-(i-1)\log(i-1)\right)$
to be a linear function. We have $\tilde{\Lambda}(X)=\Lambda(X)$. For any uniform $Y$, we have
$\tilde{\Lambda}(Y)$ $=|\mathcal{Y}|^{-1}\sum_{y\in\mathcal{Y}}\left(y\log y-(y-1)\log(y-1)\right)$
$\geq|\mathcal{Y}|^{-1}\sum_{y=1}^{|\mathcal{Y}|}\left(y\log y-(y-1)\log(y-1)\right)$
$=\log|\mathcal{Y}|.$
Hence, $\Lambda(X)=\min_{\tilde{\Lambda}}\tilde{\Lambda}(X)$.
For (7), if $X$ is conditionally uniform given $Y$, then $H(X|Y)=\Lambda(X|Y)\leq\Lambda(X)$ by concavity (the concavity of $\Lambda$ follows from (18)). To show that $H(X|Y)=\Lambda(X)$ is achievable, consider a random variable $Y\in[0,\infty)$ distributed as $Y|\{X=x\}\sim\mathrm{Unif}(0,p_{X}(x))$ (if we want $Y$ to be discrete, we can discretize $Y$ by dividing $[0,\infty)$ at the points $(p_{X}(x))_{x}$). Note that $p_{X|Y}(\cdot|y)$ is the uniform distribution over $\{x:\,p_{X}(x)\geq y\}$, and hence $H(X|Y)=\Lambda(X)$ is achievable by (5).
Before we prove (6), we first prove $H_{\infty}(p)\leq\Lambda(p)$. Let $a:=\max_{x}p_{X}(x)$. By (5) and Jensen’s inequality,
$\Lambda(p)$ $=\int_{0}^{a}|\{x:\,p(x)>t\}|\cdot\log|\{x:\,p(x)>t\}|\mathrm{d}t$
$\geq a\cdot\frac{\int_{0}^{a}|\{x:\,p(x)>t\}|\mathrm{d}t}{a}\cdot\log\frac{\int_{0}^{a}|\{x:\,p(x)>t\}|\mathrm{d}t}{a}$
$=\log(1/a)$
$=H_{\infty}(p).$
Equality holds if and only if |$|\{x:\,p(x)>t\}|$ is constant for almost all $t\in[0,a]$, meaning that $p$ is a uniform distribution.
We now prove (6). We have $H_{\infty}(X|Y)\leq\Lambda(X|Y)\leq\Lambda(X)$. Achievability of $H_{\infty}(X|Y)=\Lambda(X)$ is the same as in the proof of (7).
We finally prove (8). For every $p_{X,K}$ satisfying $p_{X,K}(x,k)\leq p_{K}(k)/k$, we have
$H_{\infty}(X|K)$ $=\mathbb{E}[-\log\max_{x}p_{X,K}(x,K)/p_{K}(K)]$
$\geq\mathbb{E}[\log K].$
Hence, $\mathbb{E}[\log K]\leq\Lambda(X)$ by (6). For the other direction, consider a random variable $Y\in[0,\infty)$ distributed as $Y|\{X=x\}\sim\mathrm{Unif}(0,p_{X}(x))$, and let $K=|\mathrm{supp}p_{X|Y}(\cdot|Y)|$. We have $p_{X,K}(x,k)\in\{0,\,p_{K}(k)/k\}$ and $\mathbb{E}[\log K]=H(X|Y)=\Lambda(X)$.
-B Proof of Proposition 3
We have already proved $H_{\infty}(X)\leq\Lambda(X)$ in Appendix -A. $\Lambda(X)\leq H(X)$ follows from (7). From direct evaluation, $\Lambda(X)=H(X)=\log|\mathcal{X}|$ for uniform $X$. To show $\Lambda(X)=H(X)$ implies that $X$ is uniform, by (7), if $\Lambda(X)=H(X)$, then there exists $Y$ such that $X$ is conditionally uniform given $Y$ and $H(X|Y)=H(X)$, implying that $Y\perp\!\!\!\perp X$, and $X$ is uniform.
The concavity of $\Lambda$ follows from (18). Schur concavity follows from concavity and the fact that $\Lambda(p)$ is invariant under labelling of $p$. Monotone linearity follows directly from the definition of $\Lambda$.
For superadditivity, let $\Lambda(X)=H(X|Z)$ and $\Lambda(Y)=H(Y|W)$ where $Z,W$ attain the maximum in (7), and $(X,Z)\perp\!\!\!\perp(Y,W)$. Since $X$ is uniform conditional on $Z$, and $Y$ is uniform conditional on $W$, $(X,Y)$ is uniform conditional on $(Z,W)$, and hence $\Lambda(X,Y)\geq H(X,Y|Z,W)=\Lambda(X)+\Lambda(Y)$ by (7). We then show the equality case. If $\Lambda(X,Y)=\Lambda(X)+\Lambda(Y)$, then $\Lambda(X,Y)=H(X,Y|Z,W)$. Let $\mathcal{X}_{z}:=\{x:\,p_{X|Z}(x|z)>0\}$, $\mathcal{Y}_{w}:=\{y:\,p_{Y|W}(y|w)>0\}$. We have $(X,Y)|\{(Z,W)=(z,w)\}\sim\mathrm{Unif}(\mathcal{X}_{z}\times\mathcal{Y}_{w})$. For $(z_{1},w_{1})\neq(z_{2},w_{2})$, if neither of $\mathcal{X}_{z_{1}}\times\mathcal{Y}_{w_{1}}$ or $\mathcal{X}_{z_{2}}\times\mathcal{Y}_{w_{2}}$ contains the other, then
$\Lambda(X,Y|(Z,W)\in\{(z_{1},w_{1}),(z_{2},w_{2})\})$
$>\sum_{i=1}^{2}\frac{p_{Z,W}(z_{i},w_{i})}{p_{Z,W}(z_{1},w_{1})+p_{Z,W}(z_{2},w_{2})}\Lambda(X,Y|(Z,W)=(z_{i},w_{i}))$
since $p_{X,Y}(x,y|z_{1},w_{1})$ has a different ordering as $p_{X,Y}(x,y|z_{2},w_{2})$, and hence $H(X,Y|Z,W)$ does not achieve the maximum of $H(X,Y|T)$ subject to that $(X,Y)$ is conditionally uniform given $Y$. Therefore, for every $(z_{1},w_{1})\neq(z_{2},w_{2})$, one of $\mathcal{X}_{z_{1}}\times\mathcal{Y}_{w_{1}}$ or $\mathcal{X}_{z_{2}}\times\mathcal{Y}_{w_{2}}$ contains the other. This is impossible if there are two possible sets for $\mathcal{X}_{z}$ and two possible sets for $\mathcal{Y}_{w}$ (in this case, we can take $\mathcal{X}_{z_{1}}\nsubseteq\mathcal{X}_{z_{2}}$ and $\mathcal{Y}_{w_{2}}\nsubseteq\mathcal{Y}_{w_{1}}$). Hence, there is only one possible set for $\mathcal{X}_{z}$ (implying $X$ is uniform), or there is only one possible set for $\mathcal{Y}_{w}$ (implying $Y$ is uniform).
For bounded increase, assume $Y\in[|\mathcal{Y}|]$, $X\in\mathbb{N}$, $p_{X}(x)$ is in descending order, and $p_{Y|X}(y|x)$ is in decreasing order with respect to $y$ for every fixed $x$ without loss of generality. Let $Z=|\mathcal{Y}|(X-1)+Y$. Then $Z$ has the same information as $(X,Y)$. Letting $\ell(i):=i\log i-(i-1)\log(i-1)$, we have
$\Lambda(X,Y)$ $=\sum_{i=1}^{\infty}p_{(X,Y)}^{\downarrow}(i)\ell(i)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{x=1}^{\infty}\sum_{y=1}^{|\mathcal{Y}|}p_{X}^{\downarrow}(x)p_{Y|X}^{\downarrow}(y|x)\ell((x-1)|\mathcal{Y}|+y)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{x=1}^{\infty}\sum_{y=1}^{|\mathcal{Y}|}p_{X}^{\downarrow}(x)\cdot|\mathcal{Y}|^{-1}\ell((x-1)|\mathcal{Y}|+y)$
$=\sum_{x=1}^{\infty}p_{X}^{\downarrow}(x)\left(x\log(x|\mathcal{Y}|)-(x-1)\log((x-1)|\mathcal{Y}|)\right)$
$=\Lambda(X)+\log|\mathcal{Y}|,$
where (a) and (b) are by rearrangement inequality where $p_{Y|X}^{\downarrow}(y|x)$ is the $y$-th largest entry of $(p_{Y|X}(y^{\prime}|x))_{y^{\prime}}$. If $\Lambda(X,Y)=\Lambda(X)+\log|\mathcal{Y}|$, then equality in (a) and (b) holds, so we must have $p_{Y|X}^{\downarrow}(y|x)=|\mathcal{Y}|^{-1}$, implying $Y$ is uniform and $X\perp\!\!\!\perp Y$.
-C Proof of Proposition 4
Assume $X\in\mathbb{N}$, and $p_{X}(x)$ is sorted in descending order. Fix $\lambda>1$. Let
$\psi(x):=c^{-1}2^{-\lambda(x\log x-(x-1)\log(x-1))}$
be a probability mass function over $\mathbb{N}$, where $c:=\sum_{x=1}^{\infty}2^{-\lambda(x\log x-(x-1)\log(x-1))}$. We have
$H(X)$ $\leq\sum_{x=1}^{\infty}p_{X}(x)\log\frac{1}{\psi(x)}$
$=\lambda\Lambda(X)+\log c.$
To bound $c$, we have
$c$ $=1+\sum_{x=2}^{\infty}2^{-\lambda(x\log x-(x-1)\log(x-1))}$
$=1+\sum_{x=2}^{\infty}2^{-\lambda\int_{x-1}^{x}\log(et)\mathrm{d}t}$
$\leq 1+\sum_{x=2}^{\infty}\int_{x-1}^{x}2^{-\lambda\log(et)}\mathrm{d}t$
$=1+\int_{1}^{\infty}(et)^{-\lambda}\mathrm{d}t$
$=1+\frac{e^{-\lambda}}{\lambda-1}$
$<1+\frac{e^{-1}}{\lambda-1}.$
The bound (9) follows from taking $\lambda=1+\eta/\Lambda(X)$.
-D Proof of Proposition 5
Without loss of generality, assume $X\in\mathbb{N}$ and $p_{X}(x)$ is in descending order. We have $L(X)=\mathbb{E}[\lfloor\log X\rfloor]$, and $\Lambda(X)=\mathbb{E}[X\log X-(X-1)\log(X-1)]$. Since $x\log x-(x-1)\log(x-1)\geq x\log x-(x-1)\log x=\log x$, we have $\mathbb{E}[\lfloor\log X\rfloor]\leq\Lambda(X)$. To show the lower bound, note that any $p_{X}$ in descending order is a convex combination of $\mathrm{Unif}([k])$ for $k\in\mathbb{N}$. Hence, it suffices to show $L(X)\geq\Lambda(X)-2$ for $X\sim\mathrm{Unif}([k])$. In this case,
$kL(X)$ $=\sum_{x=1}^{k}\lfloor\log x\rfloor$
$=\sum_{i=0}^{\lfloor\log k\rfloor-1}2^{i}\cdot i+(k-2^{\lfloor\log k\rfloor}+1)\lfloor\log k\rfloor$
$=2^{\lfloor\log k\rfloor}\left(\lfloor\log k\rfloor-2\right)+2+(k-2^{\lfloor\log k\rfloor}+1)\lfloor\log k\rfloor$
$=-2\cdot 2^{\log k-t}+2+(k+1)(\log k-t)$
$=k\log k+\log k-t(k+1)-2^{1-t}k+2,$
where $t:=\log k-\lfloor\log k\rfloor\in[0,1]$. Note that $t(k+1)+2^{1-t}k$ is a convex function in $t$, which is maximized at $t=0$ or $t=1$. Hence, $t(k+1)+2^{1-t}k\leq\max\{2k,k+1+k\}=2k+1$, and
$kL(X)$ $=k\log k+\log k-t(k+1)-2^{1-t}k+2$
$\geq k\log k+\log k-2k+1$
$>k(\Lambda(X)-2).$
-E Proof of Theorem 8
Note that $H(X\backslash Y)=H(X)$ if and only if $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$ due to the tie-breaking rule in Definition 6 (if $H(X\backslash Y)=H(X)$, then $X$ is a conditional compression, so we must choose $X\backslash Y\stackrel{{\scriptstyle\iota}}{{=}}X$ since it minimizes $H(X|U)$).
To show
$\min_{p_{Y|X}:\,H(X\backslash Y)=H(X)}H(X|Y)\leq\Lambda(X),$ (19)
consider a random variable $Y\in[0,\infty)$ distributed as $Y|\{X=x\}\sim\mathrm{Unif}(0,p_{X}(x))$ (if we want $Y$ to be discrete, we can discretize $Y$ by dividing $[0,\infty)$ at the points $(p_{X}(x))_{x}$). Note that $p_{X|Y}(\cdot|y)$ is the uniform distribution over $\{x:\,p_{X}(x)\geq y\}$, and hence $H(X|Y)=\Lambda(X)$ by (5). By Proposition 7, the probability mass function of $X\backslash Y$ is
$p_{X\backslash Y}^{\downarrow}(i)$
$=\mathbb{E}_{Y}\big{[}p_{X|Y}^{\downarrow}(i|Y)\big{]}$
$=\int_{0}^{\infty}|\{x:\,p_{X}(x)\geq y\}|\cdot p_{X|Y}^{\downarrow}(i|y)\mathrm{d}y$
$=\int_{0}^{\infty}|\{x:\,p_{X}(x)\geq y\}|\frac{\mathbf{1}\{|\{x:\,p_{X}(x)\geq y\}|\geq i\}}{|\{x:\,p_{X}(x)\geq y\}|}\mathrm{d}y$
$=\int_{0}^{\infty}\mathbf{1}\{|\{x:\,p_{X}(x)\geq y\}|\geq i\}\mathrm{d}y$
$=p_{X}^{\downarrow}(i),$ (20)
and hence $H(X\backslash Y)=H(X)$. Therefore, (19) holds.
It is left to show
$\min_{p_{Y|X}:\,H(X\backslash Y)=H(X)}H(X|Y)\geq\Lambda(X).$ (21)
Without loss of generality, assume $X\in\mathbb{N}$ and $p_{X}(x)$ is sorted in descending order. Consider any $Y$ with $H(X\backslash Y)=H(X)$. Let $U=X\backslash Y$. By Proposition 7,
$\sum_{i=1}^{k}p_{U}^{\downarrow}(i)=\mathbb{E}\Big{[}\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|Y)\Big{]}\geq\sum_{i=1}^{k}p_{X}(i),$
and hence $p_{U}\succeq p_{X}$ and $H(U)\leq H(X)$. Since equality holds ($H(U)=H(X)$), we must have $p_{U}^{\downarrow}(i)=p_{X}(i)$, and
$\mathbb{E}\Big{[}\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|Y)\Big{]}=\sum_{i=1}^{k}p_{X}(i).$
This implies $p_{X|Y}(x|y)$ must be nonincreasing with $x$ for every fixed $y$, i.e., $p_{X|Y}(x|y)=p_{X|Y}^{\downarrow}(x|y)$ (otherwise if there exists $k$ such that $\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|y)>\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|y)$, then $\mathbb{E}[\sum_{i=1}^{k}p_{X|Y}^{\downarrow}(i|Y)]>\sum_{i=1}^{k}p_{X}(i)$). By monotone linearity (Proposition 3),
$H(X|Y)$ $\geq\Lambda(X|Y)$
$=\Lambda(X).$
Therefore, (21) holds. Note that this also shows that $\Lambda(X|Y)=\Lambda(X)$ if $H(X\backslash Y)=H(X)$.
-F Proof of Theorem 11
Assume $\tilde{\Lambda}$ satisfies the conditioning property $\tilde{\Lambda}(X\backslash Y)=\tilde{\Lambda}(X|Y)$, and $\tilde{\Lambda}(X)=\log|\mathcal{X}|$ if $X$ is uniform, and $\tilde{\Lambda}(X)=\tilde{\Lambda}(Y)$ whenever $X\stackrel{{\scriptstyle\iota}}{{=}}Y$. Consider any $X$. Consider a random variable $Y\in[0,\infty)$ distributed as $Y|\{X=x\}\sim\mathrm{Unif}(0,p_{X}(x))$ (if we want $Y$ to be discrete, we can discretize $Y$ by dividing $[0,\infty)$ at the points $(p_{X}(x))_{x}$). Note that $p_{X|Y}(\cdot|y)$ is the uniform distribution over $\{x:\,p_{X}(x)\geq y\}$. We have shown in (20) that $p_{X}^{\downarrow}(i)=p_{X\backslash Y}^{\downarrow}(i)$, and hence
$\tilde{\Lambda}(X)$ $=\tilde{\Lambda}(X\backslash Y)$
$=\tilde{\Lambda}(X|Y)$
$=\mathbb{E}_{Y}\left[\tilde{\Lambda}(p_{X|Y}(\cdot|Y))\right]$
$=\mathbb{E}_{Y}\left[\tilde{\Lambda}(\mathrm{Unif}(\{x:\,p_{X}(x)\geq y\}))\right]$ (22)
$=\mathbb{E}_{Y}\left[\log|\{x:\,p_{X}(x)\geq y\}|\right]$
$=\int_{0}^{\infty}|\{x:\,p_{X}(x)\geq y\}|\log|\{x:\,p_{X}(x)\geq y\}|\mathrm{d}y$
$=\Lambda(X)$
by (5).
-G Proof of Theorem 12
Assume $\tilde{\Lambda}$ satisfies the conditioning property $\tilde{\Lambda}(X\backslash Y)=\tilde{\Lambda}(X|Y)$, $\tilde{\Lambda}(X)\leq H(X)$ for every $X$, and $\tilde{\Lambda}(X)=\tilde{\Lambda}(Y)$ whenever $X\stackrel{{\scriptstyle\iota}}{{=}}Y$. Consider any $X$, and define $Y$ as in the proof of Theorem 11. By (22),
$\tilde{\Lambda}(X)$ $=\mathbb{E}_{Y}\left[\tilde{\Lambda}(\mathrm{Unif}(\{x:\,p_{X}(x)\geq y\}))\right]$
$\leq\mathbb{E}_{Y}\left[H(\mathrm{Unif}(\{x:\,p_{X}(x)\geq y\}))\right]$
$=\mathbb{E}_{Y}\left[\log|\{x:\,p_{X}(x)\geq y\}|\right]$
$=\Lambda(X).$
-H Proof of Proposition 14
First prove that $\Lambda_{\alpha}(X)$ is non-increasing in $\alpha$. We have
$\frac{\mathrm{d}}{\mathrm{d}\beta}\Lambda_{1/\beta}(X)$
$=\frac{\mathrm{d}}{\mathrm{d}\beta}\frac{1}{\beta-1}\log\sum_{i=1}^{\infty}p^{\downarrow}(i)\big{(}i^{\beta}-(i-1)^{\beta}\big{)}$
$=\frac{\mathrm{d}}{\mathrm{d}\beta}\frac{1}{\beta-1}\log\sum_{i=1}^{\infty}i^{\beta}(p^{\downarrow}(i)-p^{\downarrow}(i+1))$
$=\frac{1}{(\beta-1)^{2}}\Bigg{(}\frac{\sum_{i=1}^{\infty}i^{\beta}(p^{\downarrow}(i)-p^{\downarrow}(i+1))\log i^{\beta-1}}{\sum_{i=1}^{\infty}i^{\beta}(p^{\downarrow}(i)-p^{\downarrow}(i+1))}$
$\quad\quad\quad\quad\quad\quad-\log\sum_{i=1}^{\infty}i^{\beta}(p^{\downarrow}(i)-p^{\downarrow}(i+1))\Bigg{)}$
$=\frac{1}{(\beta-1)^{2}}\sum_{i=1}^{\infty}p_{\beta}(i)\log\frac{p_{\beta}(i)}{q(i)}$
$=\frac{1}{(\beta-1)^{2}}D_{\mathrm{K\Lambda}}(p_{\beta}\|q)$
$\geq 0,$
where
$p_{\beta}(x):=\frac{x^{\beta}(p^{\downarrow}(x)-p^{\downarrow}(x+1))}{\sum_{i=1}^{\infty}i^{\beta}(p^{\downarrow}(i)-p^{\downarrow}(i+1))},$
$q(x):=x(p^{\downarrow}(x)-p^{\downarrow}(x+1)).$
We then prove $\Lambda_{\alpha}(X)\leq H_{\alpha}(X)$. We already have $\Lambda_{1}(X)\leq H_{1}(X)$ in Proposition 3. We first prove $\Lambda_{\alpha}(X)\leq H_{\alpha}(X)$ for $\alpha<1$. It is equivalent to
$\sum_{i=1}^{\infty}p^{\downarrow}(i)\big{(}i^{1/\alpha}-(i-1)^{1/\alpha}\big{)}\leq\left(\sum_{i=1}^{\infty}(p^{\downarrow}(i))^{\alpha}\right)^{1/\alpha}.$ (23)
Note that any non-increasing probability mass function (pmf) over $\mathbb{N}$ is a convex combination of $\mathrm{Unif}(\{1,\ldots,k\})$ for $k\geq 1$. Also, the right-hand-side of (23) is concave in $p^{\downarrow}$. Hence, it suffices to verify (23) when $p$ is the pmf of $\mathrm{Unif}(\{1,\ldots,k\})$, where (23) holds since both sides are $k^{1/\alpha-1}$. Then, we prove $\Lambda_{\alpha}(X)\leq H_{\alpha}(X)$ for $\alpha>1$. It is equivalent to
$\sum_{i=1}^{\infty}p^{\downarrow}(i)\big{(}i^{1/\alpha}-(i-1)^{1/\alpha}\big{)}\geq\left(\sum_{i=1}^{\infty}(p^{\downarrow}(i))^{\alpha}\right)^{1/\alpha}.$ (24)
The same arguments for (23) hold, except now the right hand side is convex in $p^{\downarrow}$, so the inequality is flipped. The remaining properties follow from direct computation.
-I Proof of Theorem 16
We invoke a result in [7] (also see [11, Lemma 12]): for every $X,Y$, there exists $S,K$ such that $S$ is independent of $X$, $H(K|X,S)=H(Y|K,S)=0$, and $K$ is conditionally a geometric random variable given $(X,Y)$:
$K|\{(X,Y)=(x,y)\}\sim\mathrm{Geom}(\rho(x,y)),$ (25)
where
[…]
$\geq(2^{\iota_{X;Y}(x;y)}+1)^{-1},$
where $\iota_{X;Y}(x;y):=\log\frac{\mathrm{d}P_{Y|X}(\cdot|x)}{\mathrm{d}P_{Y}}(y)$ is the information density. We have
$\Lambda(Y|S)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\Lambda(K|S)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\Lambda(K)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\Lambda(K|X,Y)$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\Lambda_{1/2}(K|X,Y)$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\mathbb{E}_{X,Y}\left[\log(2\mathbb{E}[K|X,Y]-1)\right]$
=$\stackrel{{\scriptstyle(f)}}{{=}}\mathbb{E}\left[\log(2(2^{\iota_{X;Y}(X;Y)}+1)-1)\right]$
$=\mathbb{E}\left[\log(2^{\iota_{X;Y}(X;Y)}+1/2)\right]+1,$
where (a) is because $H(Y|K,S)=0$ and the Schur concavity of $\Lambda$ (Proposition 3), (b) is by the concavity of $\Lambda$ (Proposition 3), (c) is by monotone linearity (Proposition 3) since $p_{K|X,Y}(\cdot|x,y)$ is always nonincreasing, (d) and (e) are by Proposition 14, and (f) is by (25). The remaining steps are similar to [8]:
$\Lambda(Y|S)$
$\leq\mathbb{E}\left[\log(2^{\iota_{X;Y}(X;Y)}+1/2)\right]+1$
$=\mathbb{E}\left[\log(2^{\iota_{X;Y}(X;Y)})\right]+\mathbb{E}\left[\log\left(1+2^{-\iota_{X;Y}(X;Y)-1}\right)\right]+1$
$\leq I(X;Y)+\log\left(1+\mathbb{E}\left[2^{-\iota_{X;Y}(X;Y)-1}\right]\right)+1$
$\leq I(X;Y)+\log\left(1+2^{-1}\right)+1$
$=I(X;Y)+\log 3.$
By Proposition 4, for any $\eta>0$
$H(Y|S)$
$\leq\Lambda(Y|S)+\log\left(1+\frac{\Lambda(Y|S)}{e\eta}\right)+\eta$
$\leq I(X;Y)+\log\left(1+\frac{I(X;Y)+\log 3}{e\eta}\right)+\log 3+\eta$
$=I(X;Y)+\log\left(I(X;Y)+\log 3+e\eta\right)+\log\frac{3}{e\eta}+\eta.$